# Conceptualization, Data Curation Conceptualization Visualization, Writing - Review

### v1

@@ 2 Research Methodology
This paper focuses on signal-based approaches for condition monitoring of industrial plants. As the amount of available sensor data can be abundant in real-world scenarios, the main focus of this research is on machine learning and deep learning techniques, especially neural networked-based approaches that can handle large amounts of data. For a fair comparison of the different applied methods, a simulated plant known as the Tennessee Eastman Process (TEP) is considered as the case study and different condition monitoring methods are compared based on their performance over this use case. A three-level keyword assembly is used to gather related papers and research to address the mentioned aspects. Table 1 shows the keyword breakdown of the query. All papers containing "Condition Monitoring", "Fault Detection", "Fault Diagnosis", "Prediction", "Fault Prognosis", and "Anomaly Detection" can be considered for review as they cover the application aspect of this paper. From the methodological point of view, the "Machine Learning" and "Deep Learning" approaches with a focus on "Neural Networks" have been selected. More than 300 papers are extracted from the Web of Science database using this query. The number of publications over the years is shown in Figure 2. Early papers on deep learning in this field were published in the early 2000s; however, these methods gained significant attention in recent years, especially after 2015, when research on more advanced machine learning and deep learning algorithms increased in popularity. This increase in popularity is largely attributed to improved computing resources, algorithmic breakthroughs, and the ability of such methods to handle complex big data. Within the scope of ML and DL, many different algorithms have been used for condition monitoring of industrial plants. Of the selected papers, the most frequently used algorithms are Autoencoders, Convolutional Neural Networks (CNNs), Recursive Neural Networks (RNNs), Deep Beleif Networks (DBNs), Attention mechanisms and Generative Adversarial Networks (GANs). Figure 3 shows the distribution of different algorithms among the studied papers.
Context |
Query |
Search Fields |
Application |
"Condition Monitoring" or "Fault detection" or "Fault diagnosis" or "Prediction" or "Fault Prognosis" or "Anomaly Detection" |
Title & Keywords & Abstract |
Method |
"Machine Learning" or "Deep Learning" or "Neural Network" |
Title & Keywords & Abstract |
Case Study |
"Tennessee Eastman" |
Title & Keywords & Abstract |
@@ 4.1 Recurrent Neural Network (RNN) and Long Short-Term Memory (LSTM)
Yuan et al. [113], used GRU to extract the information from correlated sequential data, a moving horizon is used at the input of the GRU to divide the raw data into the sequence units. To overcome the covariate shift and improve the network’s learning efficiency, batch normalization is used. BN-based GRU is faster and more accurate. GRU and BN-GRU accuracy is 83.42%, and 87.36% respectively.
@@ 4.2 Convolutional Neural Network (CNN)
One of the first applications of using CNN for TEP fault detection and diagnosis is introduced in [92]. In this approach, a deep convolutional neural network (DCNN) with dropout extracts features from time series data. The implemented method reached the average FDR of 88.2% for 20 faults of TEP, however, this model highly relies on historical fault data samples and is not very effective for super complex chemical processes with thousands of variables. Some modifications to the CNN network are done in [12] to increase the generalization ability of the model, improving feature extraction abilities and reducing the overfitting such as residual connections, inception module, and dilation convolution. With these improvements, the F1 score of hard detectable faults 3,9,13,15 is improved to 0.91 . Frequency domain data contain valuable information that is usually neglected in the traditional FDD of chemical processes. In [116], the Fast Fourier Transform (FFT) is used for TEP feature extraction, and the output data is fed into a convolutional neural network which is used as a binary classifier. The FDR of this model on TEP data considering all faults except 3,9,15,16 is 96.9%. In CNN all the features have the same importance in the classification process. Multi-head attention CNN (MACNN) [18] are introduced to distinguish the importance of different features, where the extracted features from convolution layers are fed into the multi-head attention (MA) layers, finally a softmax classifier is used to classify the different faults, the performance is improved compared to feedforward neural network (FNN) and CNN models. Multiblock temporal convolutional network (MBTCN) for FDD is introduced in [28] using the idea of "local extraction and global integration", it considers the cross-correlation and the temporal-correlation in the multivariate processes’ data. In this method, most related process variables are divided into the same sub-blocks where TCN extract the features. It achieved 96.2% accuracy in TEP considering all fault classes but 3,9,15. In [111], a multichannel one-dimensional convolutional neural network (MC1-DCNN) was used for FDD of the TEP process using extracted features from the Wavelet transform. Wavelet transform is a powerful tool for extracting time and frequency domain features. The average recognition accuracy of the implemented method for all 21 faults of TEP is 94.08%. To handle multi-time scale data, Gao et al. [21], introduced a feature extraction method based on CNN. The similarity measure is used to eliminate fault-irrelevant variables. The fixed multiple sampling (FMS) method is introduced for converting a feature map with an arbitrary shape into a fixed-length vector to handle different time scales of data. The average FDR for all 21 classes using the multi-time scale model is 98.53%, while the single-time scale model is 94.65%. A multi-model monitoring method is introduced in[39], in which first a CNN model detects some faults quickly and with minimum computational power, then a wavelet-assisted secondary CNN model accurately detects the remaining faults and also filters the noise and redundant information, the FDR of the model for 20 faults of TEP is 93%. A normalized CNN is used in [86] to overcome the problem of gradient explosion and disappearance, moreover the convergence speed of the network is improved. This network consists of the depth-normalized convolution network and improved second-order pooling for feature fusion. The average fault detection accuracy for all faults is 92.75%. It should be noted that this method needs a large amount of calculations. To deal with high dimensional nonlinearity the combination of matrix diagram and multi-scale convolutional neural network (MsCNN) is used in [79]. Time series data is converted into matrix diagrams (2-D images). The average fault detection accuracy for all faults is 88.54%. It should be noted that running this model is very time-consuming. Sparse one-dimensional convolutional neural network (S1-DCNN) [112], is introduced to extract features from complex signals, the method uses a sparsity regularization layer at the output of the convolution network to remove any redundant representations and extract fault-discriminative features. In [71], adaptive multiscale convolutional neural network (AMCNN), two attention modules, adaptive channel attention and adaptive spatial attention, are incorporated to select the most useful features and suppress redundant information. Another novelty in this paper is the proposed optimization method using a weighted loss function consisting of classification loss, regularization loss and triplet loss to improve robustness and generalization. the average fault detection rate for 20 faults of TEP is 93.02%. Dynamic convolutional neural networks can extract temporal correlation of process data and are useful for feature selection, and feature sequence arrangement as in [19], first, the genetic algorithm (GA) is used for feature selection and then CNN is used for fault diagnosis. The FDR for 20 faults of TEP is 89.72%. A 1D-CNN network is used in [58] for FDD. This model exploits the correlation in the data, and the average FDD accuracy for all TEP faults is 93%. Multiblock adaptive convolution kernel neural network (MBCKN) [26] is a feature extraction method for large-scale processes, in which the process is decomposed into several blocks, and the attention mechanism is used for convolutional kernel selection for each block. Finally, an adaptive fusion method fuses the features of different blocks. This model can detect 21 faults of the TEP with an average accuracy of 94.15%.
@@ 4.2.1 LSTM and CNN
Yuan et al. [114] proposed a framework for feature extraction and fault diagnosis, where the discrete wavelet transform and CNN are used for feature extraction then the LSTM network reduces the useless information and finally classification is done by using a softmax classifier. The model is validated on a few faults (Fault 1, 2, 4, 8, 10, 12, 13, 14, 17, and 18) of the TEP and the average F1-score is 0.9873. In [76], first, the LSTM network is used for feature extraction, then convolutional kernels extract characteristics of the data. This combination results in the high average fault detection accuracy of 92.06% for the TEP data set. This model can detect faults 3, 9, and 15 which are very difficult to diagnose with a very high accuracy. Bao et al. [6] used a deep convolutional neural network (DCNN) to extract spatial features from raw data and then fuse features into the bidirectional recurrent neural network (BRNN). No additional data preprocessing is needed in this case. The average accuracy for 20 faults of TEP is 91.7%. The combination of the sliding window on integrated 2D image-like data of feature information and time delay information, and the CNN-LSTM model is used in [31] for process fault detection. This method is robust against noise and converges faster compared to separate CNN and LSTM networks.
@@ 4.3 Auto-encoders
Stacked denoising autoencoder (SDAE) alongside the k-nearest neighbour (KNN) rule is used for automated feature learning in [124], the SDAE automatically extracts crucial features from nonlinear process data. Neural component analysis (NCA) for fault detection is introduced in [125], which is a unified model including a nonlinear encoder and a linear decoder, the network is trained with orthogonal constraints such as those used in PCA to alleviate the overfitting problem in autoencoders. In [15] a variational recurrent autoencoder is used for process fault detection. This method monitors the process in probability space, to handle processes nonlinearity more efficiently. Recurrent neural networks are also used to consider temporal dependency between variables. The negative variational score (NVS) is introduced as the monitoring metric, which considers reconstruction error and similarity between prior and posterior probability distributions. The model is validated on the TEP and the average fault detection rate of all faults expect 3,9,15 is 96.3%. The combination of an autoencoder and LSTM is a powerful tool for fault detection of rare events. In [61], an autoencoder is trained with offline normal data and used as an anomaly detector. The predicted faults will be classified by using an LSTM network. The average fault detection accuracy for the first 20 faults is 91.9%, and by not considering 9 and 15 faults which are very hard to detect, the average accuracy is 96.8%. The time delay of the fault detection of each class is also studied in this work. In [103], a fault detection model based on Mutual Information and Dynamic Stacked Sparse Autoencoder (SSAE) is introduced. The variables with a strong autocorrelation and their lag time of maximum autocorrelation were chosen using time series input data. Then, dynamic augmented data is fed to the SSAE for feature extraction and learning latent information in the dynamic process data. Based on selected process variables the average accuracy of fault detection of all 21 faults is 85.06%. Liu et al. [43], introduced a denoising sparse autoencoder (DSAE) to address the noise interference and nonlinearity in process variables. This method can achieve an accuracy higher than 99% for some faults. There are usually multiple operating modes in a chemical process, and fault detection models that are built based on one mode may have high false alarm rates. For multimode process monitoring and handling unseen modes a self-adaptive deep learning method based on local adaptive standardization and variational auto-encoder bidirectional long short-term memory (LAS-VB) is introduced in[93], in this research unstable tendency in the local moving window is detected. This method is not efficient for very slow-drif faults. Combining autoencoders and clustering algorithms for fault detection is a powerful tool, especially for dealing with unlabeled data. Zheng et al. [127], used a convolutional stacked autoencoder (SAE) network for feature extraction. Feature visualization is done by the t-distributed stochastic neighbour embedding (t-SNE) algorithm and DBSCAN clustering. The implemented method could detect Faults 1, 2, 4, 6, 7, 11, 13, 14, 17, 19 and 20 with average accuracy of 93%. Clustering is done unsupervised, however, a knowledge-based cluster annotation is used for labelling. Chadha et al. [11] also used a convolutional autoencoder for handling unlabeled data. The latent features are split into discriminative and reconstructive, and an auxiliary loss based on k-means clustering is used for discriminatory latent variables. The classification average accuracy for 3, 9, 15, and 21 faults which are very hard to detect is 50.04%. In [4], a source-aware autoencoder (SAAE) for detecting unseen faults and handling imbalanced data sets is introduced that uses Bidirectional long short-term memory (BiLSTM) and residual neural network (ResNet) to deal with randomness. The training can be done by using just normal data, and then fine-tuning by using emerged faulty samples. The proposed method detects unseen faults with an accuracy of 93.8%, moreover, the average accuracy of fault detection for different fault-to-normal ratios is 91.8%. In the balanced condition, the average accuracy is 94.3%. As usually there are not enough labelled samples, semi-supervised learning is developed that uses unlabeled data to help with the limited labelled data during model training. A consistency regularization autoencoder (CRAE) framework is introduced in [51], which captures correlations from labelled and unlabeled data. Previous samples are augmented with the current samples to increase the fault detection speed and accuracy. By just having the labels of just 20% of the dataset, the average fault detection accuracy of all the faults in the TE dataset except 3, 9, and 15 is 92.95%.
@@
To consider geometrical information from process signals during the feature extraction of SAEs in fault detection, manifold regularized stacked autoencoders (MRSAE) are introduced in [110] that considers the neighbourhood information. MRSAE does not need a large amount of labelled data. This model is superior in detecting certain types of faults like 5, 16, and 19, and the average FDR for 21 faults of TEP is 91.42%. To achieve interpretable analysis along the fault detection task, a novel orthogonal self-attentive variational autoencoder (OSAVA) model is introduced in [10]. Orthogonal attention extracts the correlations between variables and temporal dependency of different timesteps. The self-attention mechanisms utilized in the VSAE consider information from all timesteps to improve the performance. This model has a low detection delay an average of 24.2 minutes, the average FDR is 94.7%. To achieve effective and interpretable features for condition monitoring, a slow feature analysis-aided autoencoder (SFA-AE) is introduced in [38], which extracts deep slow variation patterns from the extracted features of AE. These slow patterns can show abnormal states. The approach uses a neural network consisting of a process variable selection (PVS) block, followed by an encoder-decoder structure and a deep slow feature analysis (DSFA) section. The PVS block utilizes 1D convolutional layers to generate an attention map that represents feature importance for selecting crucial process variables. Next, the encoder and decoder senctions use convolutional LSTM and deconvolutional LSTM layers respectively to learn spatiotemporal relationships of the input variables. The average FDR for 21 faults of TEP is 94.87%.

### v2

@@ 2 Research Methodology
This paper focuses on signal-based approaches for condition monitoring of industrial plants. As the amount of available sensor data can be abundant in real-world scenarios, the main focus of this research is on machine learning and deep learning techniques, especially neural networked-based approaches that can handle large amounts of data. For a fair comparison of the different applied methods, a simulated plant known as the Tennessee Eastman Process (TEP) is considered as the case study and different condition monitoring methods are compared based on their performance over this use case. A three-level keyword assembly is used to gather related papers and research to address the mentioned aspects. Table 1 shows the keyword breakdown of the query. All papers containing "Condition Monitoring", "Fault Detection", "Fault Diagnosis", "Prediction", "Fault Prognosis", and "Anomaly Detection" can be considered for review as they cover the application aspect of this paper. From the methodological point of view, the "Machine Learning" and "Deep Learning" approaches with a focus on "Neural Networks" have been selected. More than 400 papers are extracted from the Web of Science database using this query. The number of publications over the years is shown in Figure 2. Early papers on deep learning in this field were published in the early 2000s; however, these methods gained significant attention in recent years, especially after 2015, when research on more advanced machine learning and deep learning algorithms increased in popularity. This increase in popularity is largely attributed to improved computing resources, algorithmic breakthroughs, and the ability of such methods to handle complex big data. Within the scope of ML and DL, many different algorithms have been used for condition monitoring of industrial plants. Of the selected papers, the most frequently used algorithms are Autoencoders, Convolutional Neural Networks (CNNs), Recursive Neural Networks (RNNs), Deep Beleif Networks (DBNs), Attention mechanisms and Generative Adversarial Networks (GANs). Figure 3 shows the distribution of different algorithms among the studied papers.
Context | Query | Search Fields |
Application | "Condition Monitoring" or "Fault detection" or "Fault diagnosis" or "Prediction" or "Fault Prognosis" or "Anomaly Detection" | Title & Keywords & Abstract |
Method | "Machine Learning" or "Deep Learning" or "Neural Network" | Title & Keywords & Abstract |
Case Study | "Tennessee Eastman" | Title & Keywords & Abstract |
@@ 4.1 Recurrent Neural Network (RNN) and Long Short-Term Memory (LSTM)
Yuan et al. [117] used GRU to extract the information from correlated sequential data; a moving horizon is used at the input of the GRU to divide the raw data into the sequence units. To overcome the covariate shift and improve the network’s learning efficiency, batch normalization is used. BN-based GRU is faster and more accurate. GRU and BN-GRU accuracy is 83.42% and 87.36%, respectively. An end-to-end model fusion feature learning method (MCNN-DBiGRU) is introduced in [136], a feature-aligned multi-scale extraction model (MCNN) followed by a deep bidirectional mechanism that extracts time series features in the data, allowing the recurrent neural network to process input features both from past to future and from future to past. The experiments on all the fault samples of TEP achieve the precision of 94.11%.
@@ 4.2 Convolutional Neural Network (CNN)
One of the first applications of using CNN for TEP fault detection and diagnosis is introduced in [96]. In this approach, a deep convolutional neural network (DCNN) with dropout extracts features from time series data. The implemented method reached the average FDR of 88.2% for 20 faults of TEP, however, this model highly relies on historical fault data samples and is not very effective for super complex chemical processes with thousands of variables. Some modifications to the CNN network are done in [12] to increase the generalization ability of the model, improving feature extraction abilities and reducing the overfitting such as residual connections, inception module, and dilation convolution. With these improvements, the F1 score of hard detectable faults 3,9,13,15 is improved to 0.91 . Frequency domain data contain valuable information that is usually neglected in the traditional FDD of chemical processes. In [120], the Fast Fourier Transform (FFT) is used for TEP feature extraction, and the output data is fed into a convolutional neural network which is used as a binary classifier. The FDR of this model on TEP data considering all faults except 3,9,15,16 is 96.9%. In CNN all the features have the same importance in the classification process. Multi-head attention CNN (MACNN) [20] are introduced to distinguish the importance of different features, where the extracted features from convolution layers are fed into the multi-head attention (MA) layers, finally a softmax classifier is used to classify the different faults, the performance is improved compared to feedforward neural network (FNN) and CNN models. Multiblock temporal convolutional network (MBTCN) for FDD is introduced in [30] using the idea of "local extraction and global integration", it considers the cross-correlation and the temporal-correlation in the multivariate processes’ data. In this method, most related process variables are divided into the same sub-blocks where TCN extract the features. It achieved 96.2% accuracy in TEP considering all fault classes but 3,9,15. In [115], a multichannel one-dimensional convolutional neural network (MC1-DCNN) was used for FDD of the TEP process using extracted features from the Wavelet transform. Wavelet transform is a powerful tool for extracting time and frequency domain features. The average recognition accuracy of the implemented method for all 21 faults of TEP is 94.08%. To handle multi-time scale data, Gao et al. [23], introduced a feature extraction method based on CNN. The similarity measure is used to eliminate fault-irrelevant variables. The fixed multiple sampling (FMS) method is introduced for converting a feature map with an arbitrary shape into a fixed-length vector to handle different time scales of data. The average FDR for all 21 classes using the multi-time scale model is 98.53%, while the single-time scale model is 94.65%. A multi-model monitoring method is introduced in[41], in which first a CNN model detects some faults quickly and with minimum computational power, then a wavelet-assisted secondary CNN model accurately detects the remaining faults and also filters the noise and redundant information, the FDR of the model for 20 faults of TEP is 93%. A normalized CNN is used in [90] to overcome the problem of gradient explosion and disappearance, moreover the convergence speed of the network is improved. This network consists of the depth-normalized convolution network and improved second-order pooling for feature fusion. The average fault detection accuracy for all faults is 92.75%. It should be noted that this method needs a large amount of calculations. To deal with high dimensional nonlinearity the combination of matrix diagram and multi-scale convolutional neural network (MsCNN) is used in [82]. Time series data is converted into matrix diagrams (2-D images). The average fault detection accuracy for all faults is 88.54%. It should be noted that running this model is very time-consuming. Sparse one-dimensional convolutional neural network (S1-DCNN) [116], is introduced to extract features from complex signals, the method uses a sparsity regularization layer at the output of the convolution network to remove any redundant representations and extract fault-discriminative features. In [73], adaptive multiscale convolutional neural network (AMCNN), two attention modules, adaptive channel attention and adaptive spatial attention, are incorporated to select the most useful features and suppress redundant information. Another novelty in this paper is the proposed optimization method using a weighted loss function consisting of classification loss, regularization loss and triplet loss to improve robustness and generalization. the average fault detection rate for 20 faults of TEP is 93.02%. Dynamic convolutional neural networks can extract temporal correlation of process data and are useful for feature selection, and feature sequence arrangement as in [21], first, the genetic algorithm (GA) is used for feature selection and then CNN is used for fault diagnosis. The FDR for 20 faults of TEP is 89.72%. A 1D-CNN network is used in [60] for FDD. This model exploits the correlation in the data, and the average FDD accuracy for all TEP faults is 93%. Multiblock adaptive convolution kernel neural network (MBCKN) [28] is a feature extraction method for large-scale processes, in which the process is decomposed into several blocks, and the attention mechanism is used for convolutional kernel selection for each block. Finally, an adaptive fusion method fuses the features of different blocks. This model can detect 21 faults of the TEP with an average accuracy of 94.15%. To increase the feature extraction ability and prevent model degradation, a new multi-scale residual jagged dilated convolution neural network (MRJDCNN) model is designed in [15] by incorporating the principle of residual learning into the improved jagged dilated convolution, and then it is combined with the LSTM network delicately. This model could detect all the TEP faults with an average accuracy of 96.59%.
@@ 4.2.1 LSTM and CNN
Yuan et al. [118] proposed a framework for feature extraction and fault diagnosis, where the discrete wavelet transform and CNN are used for feature extraction then the LSTM network reduces the useless information and finally classification is done by using a softmax classifier. The model is validated on a few faults (Fault 1, 2, 4, 8, 10, 12, 13, 14, 17, and 18) of the TEP and the average F1-score is 0.9873. In [79], first, the LSTM network is used for feature extraction, then convolutional kernels extract characteristics of the data. This combination results in the high average fault detection accuracy of 92.06% for the TEP data set. This model can detect faults 3, 9, and 15 which are very difficult to diagnose with a very high accuracy. Bao et al. [6] used a deep convolutional neural network (DCNN) to extract spatial features from raw data and then fuse features into the bidirectional recurrent neural network (BRNN). No additional data preprocessing is needed in this case. The average accuracy for 20 faults of TEP is 91.7%. The combination of the sliding window on integrated 2D image-like data of feature information and time delay information and the CNN-LSTM model is used in [33] for process fault detection. This method is robust against noise and converges faster compared to separate CNN and LSTM networks. To ensure that no feature information is lost, Ren et al. [75] used a CNN to extract multi-level fault features, combine them into a unified feature space, and form hyper features through feature fusion. These hyper features are then fed into a stacked LSTM network for additional feature extraction. This approach effectively captures time series information and addresses the issue of gradient disappearance during long-term training,achieving the accuracy of 95.19% for all the faults. Zhao et al. [131], used an adaptive multiscale CNN and enhanced highway LSTM (ACEL) to tackle the problem of diverse time scales and data features in industrial settings. This method starts by automatically extracting multiscale features under varying operating conditions through a parallel convolutional module. It then applies efficient channel attention to emphasize more important channels while reducing the influence of less relevant ones. A bidirectional LSTM is used to create fine-grained hybrid features, combining contextual and local information learned by the CNN. Additionally, an enhanced highway configuration is employed to strengthen global temporal dependencies. The FDR and F1 are more than 98%.
@@ 4.3 Auto-encoders
Stacked denoising autoencoder (SDAE) alongside the k-nearest neighbour (KNN) rule is used for automated feature learning in [128], the SDAE automatically extracts crucial features from nonlinear process data. Neural component analysis (NCA) for fault detection is introduced in [129], which is a unified model including a nonlinear encoder and a linear decoder, the network is trained with orthogonal constraints such as those used in PCA to alleviate the overfitting problem in autoencoders. In [17] a variational recurrent autoencoder is used for process fault detection. This method monitors the process in probability space, to handle processes nonlinearity more efficiently. Recurrent neural networks are also used to consider temporal dependency between variables. The negative variational score (NVS) is introduced as the monitoring metric, which considers reconstruction error and similarity between prior and posterior probability distributions. The model is validated on the TEP and the average fault detection rate of all faults expect 3,9,15 is 96.3%. The combination of an autoencoder and LSTM is a powerful tool for fault detection of rare events. In [63], an autoencoder is trained with offline normal data and used as an anomaly detector. The predicted faults will be classified by using an LSTM network. The average fault detection accuracy for the first 20 faults is 91.9%, and by not considering 9 and 15 faults which are very hard to detect, the average accuracy is 96.8%. The time delay of the fault detection of each class is also studied in this work. In [107], a fault detection model based on Mutual Information and Dynamic Stacked Sparse Autoencoder (SSAE) is introduced. The variables with a strong autocorrelation and their lag time of maximum autocorrelation were chosen using time series input data. Then, dynamic augmented data is fed to the SSAE for feature extraction and learning latent information in the dynamic process data. Based on selected process variables the average accuracy of fault detection of all 21 faults is 85.06%. Liu et al. [45], introduced a denoising sparse autoencoder (DSAE) to address the noise interference and nonlinearity in process variables. This method can achieve an accuracy higher than 99% for some faults.
There are usually multiple operating modes in a chemical process, and fault detection models that are built based on one mode may have high false alarm rates. For multimode process monitoring and handling unseen modes a self-adaptive deep learning method based on local adaptive standardization and variational auto-encoder bidirectional long short-term memory (LAS-VB) is introduced in[97], in this research unstable tendency in the local moving window is detected. This method is not efficient for very slow-drif faults. Combining autoencoders and clustering algorithms for fault detection is a powerful tool, especially for dealing with unlabeled data. Zheng et al. [132], used a convolutional stacked autoencoder (SAE) network for feature extraction. Feature visualization is done by the t-distributed stochastic neighbour embedding (t-SNE) algorithm and DBSCAN clustering. The implemented method could detect Faults 1, 2, 4, 6, 7, 11, 13, 14, 17, 19 and 20 with average accuracy of 93%. Clustering is done unsupervised, however, a knowledge-based cluster annotation is used for labelling. Chadha et al. [11] also used a convolutional autoencoder for handling unlabeled data. The latent features are split into discriminative and reconstructive, and an auxiliary loss based on k-means clustering is used for discriminatory latent variables. The classification average accuracy for 3, 9, 15, and 21 faults which are very hard to detect is 50.04%. In [4], a source-aware autoencoder (SAAE) for detecting unseen faults and handling imbalanced data sets is introduced that uses Bidirectional long short-term memory (BiLSTM) and residual neural network (ResNet) to deal with randomness. The training can be done by using just normal data, and then fine-tuning by using emerged faulty samples. The proposed method detects unseen faults with an accuracy of 93.8%, moreover, the average accuracy of fault detection for different fault-to-normal ratios is 91.8%. In the balanced condition, the average accuracy is 94.3%. As usually there are not enough labelled samples, semi-supervised learning is developed that uses unlabeled data to help with the limited labelled data during model training. A consistency regularization autoencoder (CRAE) framework is introduced in [53], which captures correlations from labelled and unlabeled data. Previous samples are augmented with the current samples to increase the fault detection speed and accuracy. By just having the labels of just 20% of the dataset, the average fault detection accuracy of all the faults in the TE dataset except 3, 9, and 15 is 92.95%.
@@
To consider geometrical information from process signals during the feature extraction of SAEs in fault detection, manifold regularized stacked autoencoders (MRSAE) are introduced in [114] that considers the neighbourhood information. MRSAE does not need a large amount of labelled data. This model is superior in detecting certain types of faults like 5, 16, and 19, and the average FDR for 21 faults of TEP is 91.42%. To achieve interpretable analysis along the fault detection task, a novel orthogonal self-attentive variational autoencoder (OSAVA) model is introduced in [10]. Orthogonal attention extracts the correlations between variables and temporal dependency of different timesteps. The self-attention mechanisms utilized in the VSAE consider information from all timesteps to improve the performance. This model has a low detection delay an average of 24.2 minutes, the average FDR is 94.7%. To achieve effective and interpretable features for condition monitoring, a slow feature analysis-aided autoencoder (SFA-AE) is introduced in [40], which extracts deep slow variation patterns from the extracted features of AE. These slow patterns can show abnormal states. The approach uses a neural network consisting of a process variable selection (PVS) block, followed by an encoder-decoder structure and a deep slow feature analysis (DSFA) section. The PVS block utilizes 1D convolutional layers to generate an attention map that represents feature importance for selecting crucial process variables. Next, the encoder and decoder senctions use convolutional LSTM and deconvolutional LSTM layers respectively to learn spatiotemporal relationships of the input variables. The average FDR for 21 faults of TEP is 94.87%. To handle nonlinearity, non-Gaussian features, and dynamics in process monitoring of complex systems, independent component analysis (ICA) and adversarial autoencoder (AAE) are fused in [16].